A man weighing 690 N and a woman weighing 400 N have the same momentum. What is the ratio of the man's kinetic energy Km to that of the woman Kw
Answer #1

mm = mass of man = weight of man /9.8 = 690/9.8 = 70.4 kg

vm = velocity of man

mw = mass of woman = weight of woman /9.8 = 400/9.8 = 40.8 kg

vw = velocity of woman

Given that :

Momentum of man = momentum of woman

mm vm = mw vw

(70.4) vm = (40.8) vw

vw = 1.73 vm                             eq-1

Ratio of KE is given as

Ratio = KEm/KEw = (0.5) mm vm2 /((0.5) mw vw2 ) = (70.4/40.8) (1/(1.73)2) = 0.58